OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x^4 + 5x^3 - 2x^2 + 5x - 3; d(x) = x^2 + 1

OpenStudy (nurali):

f(x)/d(x) = (x^4 + 5x^3 - 2x^2 + 5x-3) / (x^2 + 1) you have to do this by long division: .........................x^2 + 5x - 3 ............ __________________________ x^2 + 1 ) x^4 + 5x^3 - 2x^2 + 5x - 3 ............-(x^4 + 0x^3 + x^2) <---Note this is (x^2 +1)*x^2 ......................5x^3 - 3x^2 + 5x ....................-(5x^3 +0x^2 + 5x) <---this is (x^2 +1)*5x ...............................-3x^2 + 0x - 3 .............................-(-3x^2 + 0x - 3) <and this is (x^2 +1)*(-3) ......................................… So after dividing, you end up with: x^2 + 5x -3

OpenStudy (anonymous):

This is the answer I have that is similar to yours f(x) = (x^2 + 1)( x^2 + 5x - 3)

OpenStudy (anonymous):

Is that correct?

OpenStudy (nurali):

please click best response.